# No Free Lunch from Random Feature Ensembles

### v1

@@ C.1 Sample-Bottlenecked Scaling
$E_{g}^{K}$ $\sim P^{-2\alpha\min(r,1)}$ $(\lambda\ll P^{1-\alpha})$ (C.13)
$E_{g}^{K}$ $\sim(\lambda/P)^{2\min(r,1)}$ $(\lambda\gg P^{1-\alpha})$ (C.14)
@@
C.2 Overparameterized Ensembles with Small Ridge
@@
$\kappa_{2}$ $\approx\kappa_{2}^{*}+\delta/N+\alpha\lambda$ (C.15)
$\operatorname{Df}_{1}(\kappa_{2})$ $\approx P+\operatorname{Df}_{1}^{\prime}(\kappa_{2}^{*})(\delta/N+\alpha\lambda)$ (C.16)
$\operatorname{Df}_{2}(\kappa_{2})$ $\approx\operatorname{Df}_{2}(\kappa_{2}^{*})+\operatorname{Df}_{2}^{\prime}(\kappa_{2}^{*})(\delta/N+\alpha\lambda),$ (C.17)
@@
$\alpha$ $=-\frac{1}{\kappa_{2}^{*}\operatorname{Df}_{1}^{\prime}(\kappa_{2}^{*})}$ (C.18)
$\delta$ $=\frac{\lambda P\kappa_{2}^{*}}{2\lambda-{\kappa_{2}^{*}}^{2}\operatorname{Df}_{1}^{\prime}(\kappa_{2}^{*})}$ (C.19)
@@
$\operatorname{Bias}_{z}^{2}$ $=-\frac{P{\kappa_{2}^{*}}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2}^{*})}{P-\operatorname{Df}_{2}(\kappa_{2}^{*})}+\lambda F(\kappa_{2}^{*},P)+\mathcal{O}(\lambda^{2},\lambda/N,1/N^{2})$ (C.20)
$\operatorname{Var}_{z}$ $=\frac{P\kappa_{2}^{*}\operatorname{tf}_{1}(\kappa_{2}^{*})}{N}+\mathcal{O}(\lambda^{2},\lambda/N,1/N^{2})\,,$ (C.21)
@@
$F(\kappa_{2}^{*},P)$ $\equiv\frac{P\left((P-\operatorname{Df}_{2}(\kappa_{2}^{*}))\operatorname{tf}_{2}^{\prime}(\kappa_{2}^{*})+\kappa_{2}^{*}\operatorname{Df}_{2}^{\prime}(\kappa_{2}^{*})\operatorname{tf}_{1}^{\prime}(\kappa_{2}^{*})\right)}{\operatorname{Df}_{1}^{\prime}(\kappa_{2}^{*})(P-\operatorname{Df}_{2}(\kappa_{2}^{*}))^{2}}$ (C.22)
@@
$E_{g}^{K}=-\frac{P{\kappa_{2}^{*}}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2}^{*})}{P-\operatorname{Df}_{2}(\kappa_{2}^{*})}+\lambda F(\kappa_{2}^{*},P)+\frac{P\kappa_{2}^{*}\operatorname{tf}_{1}(\kappa_{2}^{*})}{KN}+\mathcal{O}(\lambda^{2},\lambda/N,1/N^{2})$ (C.23)
@@
Appendix D RFRR on Real Datasets
D.1 Numerical Experiments with Synthetic Gaussian Data
@@
For a given value of $\alpha$ and $r$, we fix a large value $D\gg P,M$ and generate eigenvalues $\eta_{t}\propto t^{-\alpha}$ and target weights $\bar{w}_{t}\sim t^{-\frac{1}{2}(1-\alpha+2\alpha r)}$. The $\eta_{t}$ are normalized so that $\sum_{t}\eta_{t}=1$ and $\sum_{t}\bar{w}_{t}^{2}\eta_{t}=1$. We generate random features as $\bm{\Theta}=\bm{\Sigma}\bm{X}\in\mathbb{R}^{D\times P}$, where $\bm{X}_{ij}\sim\mathcal{N}(0,1)$ and $\bm{\Sigma}$ is the diagonal matrix with entry $\bm{\Sigma}_{tt}=\eta_{t}$. Labels are assigned as $\bm{y}=\bm{\Psi}^{\top}\bar{\bm{w}}\in\mathbb{R}^{P}$. For each $k=1,\dots,K$, we perform linear RFRR (eq. A.12) with an independently drawn projection matrix $\bm{Z}^{k}$ with entries drawn from $\mathcal{N}(0,1/N)$. The prediction of the ensemble is then given as the mean over the $K$ learned predictors.
@@
D.2 Numerical Experiments on Binarized MNIST and CIFAR-10 with ReLU Features
@@
We perform RFRR on CIFAR-10 and MNIST datasets. To construct the dataset, we sort the images into two classes. For CIFAR-10, we assign a label $y=+1$ to images of “things one could ride” together (airplane, automobile, horse, ship, truck) and a label $y=-1$ for “things one ought not to ride” (bird, cat, deer, dog, frog) (Simon et al., 2023). For MNIST, we assign a label $y=+1$ to digits $0-4$, and a label $-1$ to digits $5-9$. We construct $K$ feature maps as $\psi(^{k}\bm{x})=\frac{1}{\sqrt{N}}\operatorname{ReLU}\left(\bm{V}^{k\top}\bm{x}\right)$, where $\bm{V}^{k}_{ij}\sim\mathcal{N}(0,2/D)$. Here, $D$ is the data dimensionality ($D=3072$ for CIFAR-10 and $784$ for MNIST). Then, for each ensemble member $k=1,\dots,K$, we train a linear regression model on the features $\psi^{k}$. In the infinite-feature limit, the finite-feature kernels will converge to the “NNGP kernel” for a single-hidden-layer Relu network (Lee et al., 2018).
D.3 Theoretical Predictions
@@
We evaluate the omniscient risk estimate 16 numerically using vectors storing the values of $\{\eta_{t}\}_{t=1}^{\infty}$ and $\{\bar{w}_{t}\}_{t=1}^{\infty}$. In the case of synthetic data, these vectors are readily available. For the MNIST and CIFAR-10 tasks, we approximate these vectors by evaluating the infinite-width neural network gaussian process (NNGP) kernel using the neural tangents library (Novak et al., 2019). For 30,000 images from the training sets of both MNIST and CIFAR 10, we evaluate the kernel matrix $[\bm{H}]_{p,p^{\prime}}=\bm{H}(\bm{x}_{p},\bm{x}_{p^{\prime}})$, and diagonalize the kernel matrix to determine the eigenvectors and eigenvalues. To be precise, with $P=30,000$, we calculate the eigenvalues $\{\eta_{1},\eta_{2},\dots,\eta_{P}\}$ and eigenvectors $\{\bm{u}_{1},\dots,\bm{u}_{P}\}$ of the sample-normalized kernel matrix $\frac{1}{P}\bm{H}$. We then assign the weights of the target function as $\bar{w}_{t}=\frac{1}{\sqrt{P\eta_{t}}}\bm{u}_{t}^{\top}\bm{y}$, where $\bm{y}\in\mathbb{R}^{P}$ is the vector of labels associated to our $P$ samples.
@@
D.4 Measuring Power Law Exponents of Kernel Ridge Regression Tasks
In fig. 4B, we plot theoretical predictions for the scaling exponents of the bias and variance contributions to error for the binarized CIFAR-10 and MNIST classification tasks. To calculate these exponents, we need access to the “ground truth” source and capacity exponents $\hat{\alpha}$ and $\hat{r}$ characterizing the kernel eigenstructure of the dataset and the target function, such that the eigenvalues $\{\eta_{1},\eta_{2},\dots\}$ of the NNGP kernel decay as $\eta_{t}\sim t^{-\hat{\alpha}}$ and the weights of the target function $\{\bar{w}_{1},\bar{w}_{2},\dots\}$ decay as $\eta_{t}\bar{w}_{t}^{2}\sim t^{-(1+2\hat{\alpha}\hat{r})}$. To estimate $\hat{\alpha}$, we calculate the “trace metric” $\left[\operatorname{tr}\left[\bm{H}_{p}^{-1}\right]\right]^{-1}$ $\hat{\alpha}$ is then obtained by fitting to the relationship $\left[\operatorname{tr}\left[\bm{H}_{p}^{-1}\right]\right]^{-1}\sim p^{-\alpha}$, where $\bm{H}_{p}\in\mathbb{R}^{p\times p}$ is the empirical NNGP kernel for $p$ randomly drawn samples from the dataset (Wei et al., 2022; Simon et al., 2023) (figs. S7.A,C). To estimate the source exponent $r$, we use the scaling law for Kernel Ridge Regression (with the limiting infinite-feature NNGP kernel) which dictates that for small ridge, $E_{g}\sim p^{-2\alpha\min(r,1)}$ (see Appendix C.1). Following Simon et al., we fit the MSE loss for Kernel Ridge Regression with the limiting NNGP kernel to a power law decay $E_{g}\sim p^{-\beta}$ (figs. S7.B,D), and assign $\hat{r}=\beta/2/\hat{\alpha}$.
Appendix E Deep Ensemble Experimental Details
E.1 Natural Language Processing Tasks
The transformers used for the WikiText2 experiments are implemented in the maximal update parameterization following the setup in (Bordelon et al., 2024c). We fix depth $L=6$, number of heads $H=12$, and vary width $N$ such that $k\sqrt{N}\approx 60$ to keep the parameter count between ensembles approximately the same since parameters scale quadratically with width. The transformer is composed of alternating blocks of causal multiheaded attention and a 2-layer MLP. We make sure to center the models’ outputs and vary feature learning in the models by dividing the output by $\gamma=\gamma_{0}\sqrt{N}$.
E.2 Image Classification Tasks
The CNNs used for the CIFAR-10 experiments consist of two CNN layers and an MLP layer. To compare ensembles of different sizes fairly while keeping the total number of parameters fixed, we adopt a fixed ratio between layer sizes in each CNN. If the first CNN layer has width (channels) $c$, then the second CNN layer has width $2c$, and the MLP has width $5c$. The value of $c$ is varied such that the total ensemble maintains approximately the same number of parameters $M$ across varying numbers of ensemble members $K$. There are about $M=350$ thousand total parameters in these models.
We also examine a standard ResNet18 architecture, varying the initial width (number of output channels) in each block to maintain the same number of parameters across ensembles of varying size. There are about $M=30$ million total parameters in these models.
We employ standard data augmentation techniques during training. All training images are randomly cropped, resized, and flipped horizontally. Our models are trained using SGD with base learning rate $0.1$, momentum $0.9$, and early stopping, with a patience of five epochs.

### v2

@@ D.1 Sample-Bottlenecked Scaling
$E_{g}^{K}$ $\sim P^{-2\alpha\min(r,1)}$ $(\lambda\ll P^{1-\alpha})$ (D.13)
$E_{g}^{K}$ $\sim(\lambda/P)^{2\min(r,1)}$ $(\lambda\gg P^{1-\alpha})$ (D.14)
@@
D.2 Overparameterized Ensembles with Small Ridge
@@
$\kappa_{2}$ $\approx\kappa_{2}^{*}+\delta/N+\alpha\lambda$ (D.15)
$\operatorname{Df}_{1}(\kappa_{2})$ $\approx P+\operatorname{Df}_{1}^{\prime}(\kappa_{2}^{*})(\delta/N+\alpha\lambda)$ (D.16)
$\operatorname{Df}_{2}(\kappa_{2})$ $\approx\operatorname{Df}_{2}(\kappa_{2}^{*})+\operatorname{Df}_{2}^{\prime}(\kappa_{2}^{*})(\delta/N+\alpha\lambda),$ (D.17)
@@
$\alpha$ $=-\frac{1}{\kappa_{2}^{*}\operatorname{Df}_{1}^{\prime}(\kappa_{2}^{*})}$ (D.18)
$\delta$ $=\frac{\lambda P\kappa_{2}^{*}}{2\lambda-{\kappa_{2}^{*}}^{2}\operatorname{Df}_{1}^{\prime}(\kappa_{2}^{*})}$ (D.19)
@@
$\operatorname{Bias}_{z}^{2}$ $=-\frac{P{\kappa_{2}^{*}}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2}^{*})}{P-\operatorname{Df}_{2}(\kappa_{2}^{*})}+\lambda F(\kappa_{2}^{*},P)+\mathcal{O}(\lambda^{2},\lambda/N,1/N^{2})$ (D.20)
$\operatorname{Var}_{z}$ $=\frac{P\kappa_{2}^{*}\operatorname{tf}_{1}(\kappa_{2}^{*})}{N}+\mathcal{O}(\lambda^{2},\lambda/N,1/N^{2})\,$, (D.21)
@@
$F(\kappa_{2}^{*},P)$ $\equiv\frac{P\left((P-\operatorname{Df}_{2}(\kappa_{2}^{*}))\operatorname{tf}_{2}^{\prime}(\kappa_{2}^{*})+\kappa_{2}^{*}\operatorname{Df}_{2}^{\prime}(\kappa_{2}^{*})\operatorname{tf}_{1}^{\prime}(\kappa_{2}^{*})\right)}{\operatorname{Df}_{1}^{\prime}(\kappa_{2}^{*})(P-\operatorname{Df}_{2}(\kappa_{2}^{*}))^{2}}$ (D.22)
@@
$E_{g}^{K}=-\frac{P{\kappa_{2}^{*}}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2}^{*})}{P-\operatorname{Df}_{2}(\kappa_{2}^{*})}+\lambda F(\kappa_{2}^{*},P)+\frac{P\kappa_{2}^{*}\operatorname{tf}_{1}(\kappa_{2}^{*})}{KN}+\mathcal{O}(\lambda^{2},\lambda/N,1/N^{2})$ (D.23)
@@
Appendix E RFRR on Real Datasets
E.1 Numerical Experiments with Synthetic Gaussian Data
@@
For a given value of $\alpha$ and $r$, we fix a large value $D\gg P,M$ and generate eigenvalues $\eta_{t}\propto t^{-\alpha}$ and target weights $\bar{w}_{t}\sim t^{-\frac{1}{2}(1-\alpha+2\alpha r)}$. The $\eta_{t}$ and $\bar{w}_{t}$ are normalized so that $\sum_{t}\eta_{t}=1$ and $\sum_{t}\bar{w}_{t}^{2}\eta_{t}=1$. We generate random features as $\bm{\Theta}=\bm{\Sigma}\bm{X}\in\mathbb{R}^{D\times P}$, where $\bm{X}_{ij}\sim\mathcal{N}(0,1)$ and $\bm{\Sigma}$ is the diagonal matrix with entry $\bm{\Sigma}_{tt}=\eta_{t}$. Labels are assigned as $\bm{y}=\bm{\Psi}^{\top}\bar{\bm{w}}\in\mathbb{R}^{P}$. For each $k=1,\dots,K$, we perform linear RFRR (eq. A.12) with an independently drawn projection matrix $\bm{Z}^{k}$ with entries drawn from $\mathcal{N}(0,1/N)$. The prediction of the ensemble is then given as the mean over the $K$ learned predictors. We use this procedure to generate numerical error curves in Figures 4 and S2.A-C.
@@
E.2 Numerical Experiments on Binarized MNIST and CIFAR-10 with ReLU Features
@@
We perform RFRR on CIFAR-10 and MNIST datasets. To construct the dataset, we sort the images into two classes. For CIFAR-10, we assign a label $y=+1$ to images of “things one could ride” together (airplane, automobile, horse, ship, truck) and a label y=1y=-1 for “things one ought not to ride” (bird, cat, deer, dog, frog) (Simon et al., 2023). For MNIST, we assign a label $y=+1$ to digits 040-4, and a label 1-1 to digits 595-9. We construct $K$ feature maps as $\psi(^{k}\bm{x})=\frac{1}{\sqrt{N}}\operatorname{ReLU}\left(\bm{V}^{k\top}\bm{x}\right)$, where $\bm{V}^{k}_{ij}\sim\mathcal{N}(0,2/D)$. Here, $D$ is the data dimensionality ($D=3072$ for CIFAR-10 and $784$ for MNIST). Then, for each ensemble member $k=1,\dots,K$, we train a linear regression model on the features $\psi^{k}$. In the infinite-feature limit, the finite-feature kernels will converge to the “NNGP kernel” for a single-hidden-layer Relu network (Lee et al., 2018). This procedure was used to generate numerical error curves for figures 1, 2, 5, S1, S2.D-F, S3, S4, S5, S6, and S7
E.3 Theoretical Predictions
@@
We evaluate the omniscient risk estimate 20 numerically using vectors storing the values of $\{\eta_{t}\}_{t=1}^{\infty}$ and $\{\bar{w}_{t}\}_{t=1}^{\infty}$. In the case of synthetic data, these vectors are readily available. For the MNIST and CIFAR-10 tasks, we approximate these vectors by evaluating the infinite-width Neural Network Gaussian Process (NNGP) kernel using the neural tangents library (Novak et al., 2019). For 30,000 images from the training sets of both MNIST and CIFAR 10, we evaluate the kernel matrix $\bm{H}]_{p,p^{\prime}}=\bm{H}(\bm{x}_{p},\bm{x}_{p^{\prime}})$, and diagonalize the kernel matrix to determine the eigenvectors and eigenvalues. To be precise, with $P=30,000$, we calculate the eigenvalues $\{\eta_{1},\eta_{2},\dots,\eta_{P}\}$ and eigenvectors $\{\bm{u}_{1},\dots,\bm{u}_{P}\}$ of the sample-normalized kernel matrix $\frac{1}{P}\bm{H}$. We then assign the weights of the target function as $\bar{w}_{t}=\frac{1}{\sqrt{P\eta_{t}}}\bm{u}_{t}^{\top}\bm{y}$, where $\bm{y}\in\mathbb{R}^{P}$ is the vector of labels associated to our $P$ samples.
@@
E.4 Measuring Power Law Exponents of Kernel Ridge Regression Tasks
In fig. 5B, we plot theoretical predictions for the scaling exponents of the bias and variance contributions to error for the binarized CIFAR-10 and MNIST classification tasks. To calculate these exponents, we need access to the “ground truth” source and capacity exponents $\hat{\alpha}$ and $\hat{r}$ characterizing the kernel eigenstructure of the dataset and the target function, such that the eigenvalues $\{\eta_{1},\eta_{2},\dots\}$ of the NNGP kernel decay as $\eta_{t}\sim t^{-\hat{\alpha}}$ and the weights of the target function $\{\bar{w}_{1},\bar{w}_{2},\dots\}$ decay as $\eta_{t}\bar{w}_{t}^{2}\sim t^{-(1+2\hat{\alpha}\hat{r})}$. To estimate $\hat{\alpha}$, we calculate the “trace metric” $\left[\operatorname{tr}\left[\bm{H}_{p}^{-1}\right]\right]^{-1}$ $\hat{\alpha}$ is then obtained by fitting to the relationship $\left[\operatorname{tr}\left[\bm{H}_{p}^{-1}\right]\right]^{-1}\sim p^{-\alpha}$, where $\bm{H}_{p}\in\mathbb{R}^{p\times p}$ is the empirical NNGP kernel for $p$ randomly drawn samples from the dataset (Wei et al., 2022; Simon et al., 2023) (figs. S7.A,C). To estimate the source exponent $r$, we use the scaling law for Kernel Ridge Regression (with the limiting infinite-feature NNGP kernel) which dictates that for small ridge, $E_{g}\sim p^{-2\alpha\min(r,1)}$ (see Appendix D.1). Following Simon et al., we fit the MSE loss for Kernel Ridge Regression with the limiting NNGP kernel to a power law decay $E_{g}\sim p^{-\beta}$ (figs. S7.B,D), and assign $\hat{r}=\beta/2/\hat{\alpha}$.
Appendix F Code Availability
All code used to generate the figures presented in this work is publicly available at https://github.com/benruben87/Random_Feature_Ensembles.git.